OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5). help please i always get point slope form mixed up with y=mx=b im not sure which one to use?

OpenStudy (anonymous):

use \[y-y_1=m(x-x_1)\] with \[y_1=5,x_1=-2, m=-1\]

OpenStudy (anonymous):

you have it right, y=mx+b is point intercept form, the m is -1, so plug into the equation the values you have for x, y, and m to find b

OpenStudy (anonymous):

y=mx+b 5=-1(-2)=b 5=2+b 3=b so the equation is y=-1x+3

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

y= mx+b is slope-intercept form, not point-slope form